If f(x) = -4x +5 and g(x) = |x-6| then what is f(g)(-2))
Answer by stanbon(75887)

If f(x) = -4x +5 and g(x) = |x-6| then what is f(g)(-2))
------
f[g(-2)] = f[-2-6|] = f(8) = -4*8+5 = -27
============================
Cheers,
Stan H.
=======